# Retrieval Pivot Attacks in Hybrid RAG: Measuring and Mitigating Amplified Leakage from Vector Seeds to Graph Expansion

### v1

@@ VII-D Connectivity Sensitivity
VII-E Embedding Model Sensitivity
To verify that the pivot vulnerability is structural rather than embedding-dependent, we repeat the P1 and P3 evaluations using all-mpnet-base-v2 (768 dimensions, higher retrieval quality) in addition to our default all-MiniLM-L6-v2 (384 dimensions). We rebuild the ChromaDB collection for each model and run the full 500-query evaluation (Table VI).
TABLE VI: Embedding model sensitivity: P3 RPR and mean Leakage@k across two embedding models. The vulnerability persists regardless of embedding quality.
@@ VII-E Embedding Model Sensitivity
VII-F Attack Evaluation
Table VII presents RPR under each attack across pipeline variants. All four attacks achieve $\text{RPR}=1.0$ against the undefended hybrid pipeline (P3)—every adversarial query produces cross-tenant leakage. Critically, all four attacks achieve $\text{RPR}=0.0$ against every defended variant (P4, P6, P8).
TABLE VII: RPR under each attack type across pipeline variants. Mean Leakage@k shown in parentheses. All attacks fail against D1-defended pipelines.
@@ VII-F Attack Evaluation
Table VIII provides injection details for each attack. The attacks span a range of injection budgets (9–20 chunks) and entity strategies (1–2 target entities, 1–6 MENTIONS edges created).
TABLE VIII: Attack injection details. All attacks achieve RPR=1.0 on P3 and RPR=0.0 on P4/P6/P8.
@@
VII-G Adaptive Attacks (A5–A7)
We extend the attack taxonomy with three adaptive strategies that target the defense mechanisms themselves (Table IX).
TABLE IX: RPR under adaptive attacks (A5–A7) across pipeline variants. A5 tests metadata forgery at three rates; A6 tests entity manipulation; A7 tests query manipulation. Leakage@k shown in parentheses.
@@ VII-G Adaptive Attacks (A5–A7)
VII-H Generation Impact
To measure whether leaked context actually contaminates LLM-generated answers, we evaluate three production LLMs on 10 queries with known leakage (Table X).
TABLE X: Generation contamination metrics. ECR (Entity Contamination Rate) measures leaked entity names in answers. ILS (Information Leakage Score) measures semantic similarity to leaked content. FCR (Factual Contamination Rate) and GRR (Generation Refusal Rate) assess factual leakage and model self-censoring.
LLM ECR ILS FCR GRR Cost
GPT-4o 0.077 0.305 0.050 0.800 \$0.08
Claude Sonnet 4.5 0.307 0.346 0.065 0.100 \$0.18
DeepSeek-V3 0.124 0.336 0.190 0.200 \$0.01
Note: 10 queries with leakage evaluated per provider on synthetic dataset. Judge model: same provider as generator for each row. GPT-4o exhibits high refusal rate (GRR = 0.80), largely ignoring leaked context.
The results reveal model-dependent contamination behavior. Claude Sonnet 4.5 exhibits the highest Entity Contamination Rate ($\text{ECR}=0.31$) and lowest Generation Refusal Rate ($\text{GRR}=0.10$), indicating it readily incorporates leaked entities into generated answers. GPT-4o shows the opposite pattern: $\text{ECR}=0.08$ and $\text{GRR}=0.80$, largely ignoring leaked context. DeepSeek-V3 falls between ($\text{ECR}=0.12$, $\text{GRR}=0.20$) but produces the highest Factual Contamination Rate ($\text{FCR}=0.19$), meaning it synthesizes leaked facts into answers more than the other models. These findings confirm that retrieval-level leakage translates to generation-level contamination for most LLMs, and that the severity depends on the model’s tendency to incorporate all context versus selectively ignoring irrelevant content.
VII-I Traversal Parameter Sweep
@@
VII-J Latency and Overhead
Table XI presents latency and context size across pipeline variants.
TABLE XI: Latency (ms) and context size across pipeline variants.
@@ VII-J Latency and Overhead
VII-K Metadata Integrity Stress Test
@@ VII-M Utility Impact
Table XII presents the security-utility tradeoff. P3 (undefended hybrid) returns 110 items per query, but only 85.5% are authorized—the remaining 16 items (14.5%) are cross-tenant or over-clearance leakage. P1 (vector-only) returns 10 fully authorized items, establishing the vector-only utility baseline.
@@
VII-N Cross-Dataset Validation
To verify that Retrieval Pivot Risk is not an artifact of synthetic corpus construction, we repeat the baseline evaluation (P1, P3, P4) on the Enron email corpus described in §VI-C. Table XIII compares results across both corpora.
TABLE XIII: Cross-dataset baseline: retrieval security metrics for undefended (P3) and defended (P4) hybrid RAG across synthetic and Enron corpora. RPR = Retrieval Pivot Risk, Leak = mean leaked items per query, PD = Pivot Depth (hops), Ctx = mean context size.
@@ VII-N Cross-Dataset Validation
Note: Synthetic corpus: 1000 documents, 4 tenants, 40 bridge entities. Enron corpus: 50,000 emails, 5 departments, 19 cross-department entities. Both corpora evaluated with 200 benign queries.
Three findings emerge from the cross-dataset comparison:
The vulnerability generalizes. Enron’s undefended hybrid pipeline (P3) produces $\text{RPR}=0.695$ on benign queries—139 of 200 queries leak cross-department content, with a mean of 7.1 leaked items per query. Under adversarial queries, $\text{RPR}=0.615$ (123 of 200 queries leak). The vulnerability is not an artifact of synthetic construction: real corporate email with natural departmental boundaries exhibits the same pivot mechanism.
RPR scales with entity connectivity. The Enron corpus produces lower RPR (0.695 vs. 0.954) and lower mean leakage (7.1 vs. 16.0) than the synthetic corpus. This difference is explained by entity connectivity: the synthetic graph has 40 intentional bridge entities across 4 tenants with curated cross-references, while Enron has 19 naturally shared entities across 5 departments with sparser cross-department email traffic. The Enron graph is $135\times$ larger (376K vs. 2.8K nodes) but its entity mentions are noisier and more fragmented, reducing the density of cross-tenant pivot paths. This confirms that RPR is a function of entity connectivity density, not corpus size.
D1 eliminates all leakage on both corpora. P4 achieves $\text{RPR}=0.0$ on Enron with context size 25 (vs. 56 on the synthetic corpus). The smaller post-D1 context reflects Enron’s higher entity-to-chunk ratio—more entity nodes are filtered, leaving fewer authorized chunk traversals. The defense generalizes without modification: the same per-hop tenant-and-sensitivity check eliminates leakage regardless of corpus size, entity distribution, or graph topology.
PD = 2 persists. All leakage in both corpora occurs at exactly $\text{PD}=2$, confirming that the bipartite pivot (chunk $\to$ entity $\to$ chunk) is a structural invariant of hybrid RAG systems that construct knowledge graphs via entity linking, independent of the underlying document domain.
VIII Mitigations
We propose five defenses that operate at different stages of the hybrid pipeline, forming a defense-in-depth architecture (Table XIV).
TABLE XIV: Defense mechanisms, pipeline integration points, and experimental impact.
@@ VIII-A D1: Per-Hop Authorization
This entity-level filtering is precisely the mechanism by which D1 eliminates cross-tenant leakage. The 2-hop pivot path (chunk $\to$ entity $\to$ chunk) requires traversal through a shared entity node. By filtering entity nodes, D1 severs this path at hop 1, preventing the traversal from reaching the unauthorized chunk at hop 2. The trade-off is that entity information (which may be useful for answer generation) is excluded from the final context. We quantify this impact in Section IX.
@@
VIII-B D1 Metadata Integrity Assumption
D1 assumes that node metadata (tenant, sensitivity) is trustworthy. If an attacker can forge metadata during document injection, D1 can be bypassed. Our metadata stress test (§VII-K) shows that D1 is robust to random mislabeling up to 5%, but targeted metadata forgery is a different threat. Mitigations include: (a) system-assigned metadata during ingestion (the uploader cannot choose tenant or sensitivity labels), (b) cryptographic signing of metadata at ingestion time, and (c) periodic metadata audits that compare node labels against source-of-truth identity systems. In practice, most enterprise document management systems already enforce system-assigned tenant labels, making (a) the natural deployment model.
VIII-C D2: Edge Allowlist
@@
VIII-D D3: Budgeted Traversal
Budgeted traversal enforces hard limits on BFS expansion: maximum hop depth, maximum branching factor per node, and maximum total expanded nodes. Our traversal sweep (§VII-I) shows that $\text{total\_nodes}\leq 25$ eliminates leakage even without D1, while the default budget ($d_{\text{max}}=2$, branching $\leq 10$, total $\leq 50$) reduces context from 56 items (D1 only) to 28 items, removing authorized but irrelevant graph content that adds noise.
VIII-E D4: Trust-Weighted Expansion
@@
VIII-F D5: Merge-Time Policy Filter
@@
IX Discussion
IX-A Security-Utility Tradeoff
@@
D1’s security guarantee comes at a specific cost: entity nodes are excluded from the final context. In our bipartite construction, the undefended P3 pipeline returns approximately 110 items per query, of which roughly 25–35 are entity nodes. After D1 filtering, these entity nodes are removed entirely, leaving 50–56 chunk-only items. The entity information they carried (names, relationships, organizational context) is lost. For queries that benefit from entity-level knowledge (e.g., “Who works on project Alpha?”), this may reduce answer quality compared to an oracle that retains only authorized entity nodes. We consider this an acceptable trade-off: the entity nodes are exactly the pivot vectors that enable cross-tenant leakage, and their removal is the mechanism by which D1 achieves $\text{RPR}=0.0$. Future work on multi-tenant entity authorization (Section X) could recover entity utility while preserving security.
@@
IX-B Why PD = 2 Across Both Corpora
@@
The Enron corpus confirms this structural invariant: despite a $135\times$ larger graph (376K vs. 2.8K nodes), noisier NER extractions, and real-world email text rather than curated templates, all leakage occurs at exactly $\text{PD}=2$. The synthetic corpus has 40 shared entities; Enron has 19. Both produce the same pivot depth. This confirms that $\text{PD}=2$ is a consequence of the bipartite construction, not of corpus-specific properties.
@@ IX-C The Amplification Mechanics
IX-D Comparison with Prior Art
@@
GRAGPoison [8] demonstrates that graph structure amplifies small poisoning inputs (98% ASR from relation injection). Our work extends this insight to hybrid architectures: the combination of vector retrieval and graph expansion creates a compound attack surface where authorized vector results seed unauthorized graph traversals.
@@
PoisonedRAG [28] achieves 90% ASR by injecting 5 documents into vector stores. In vector-only RAG, this is contained by prefilters. In hybrid RAG, even a single retrieved document that mentions a shared entity can trigger graph expansion into sensitive neighborhoods, making the effective attack surface orders of magnitude larger.
@@
The Graph RAG Privacy Paradox [10] observes that graph RAG reduces text leakage but increases structural information extraction. Our findings align: the hybrid pipeline’s graph expansion systematically transforms authorized text into unauthorized structural access.
IX-E Implications for Agentic Systems
@@
The pivot attack is especially dangerous in agentic RAG deployments (LangGraph, CrewAI) where graph traversal is performed autonomously by an LLM agent [6]. In these systems, the agent decides traversal depth, edge types, and expansion strategies without human oversight. A poisoned vector seed can manipulate the agent into performing unrestricted graph exploration, and research shows that a single compromised agent can poison 87% of downstream decision-making within 4 hours [6]. Per-hop authorization is even more critical in agentic settings, where there is no human in the loop at the expansion step.
X Limitations
@@
Two corpora. We evaluate on a synthetic enterprise corpus and the Enron email corpus. The synthetic corpus validates the vulnerability under controlled conditions; the Enron corpus confirms generalization to real-world data. However, both use a bipartite chunk–entity graph topology that produces a uniform PD=2 signature. Knowledge graphs with richer entity-to-entity relationships (ontological hierarchies, multi-hop inference chains) may exhibit leakage at deeper pivot depths. We do not evaluate on production enterprise graphs with RBAC/ABAC policies, which may exhibit different connectivity patterns.
@@
Limited generation evaluation. Our generation contamination experiment (§VII-H) evaluates 10 queries per LLM on the synthetic corpus and shows that leaked context does affect generated answers (ECR up to 0.31, FCR up to 0.19). However, the sample size is small and limited to one corpus. A larger-scale generation study across both corpora would strengthen the end-to-end impact claim. We also do not evaluate chain-of-thought reasoning, which might amplify the impact by synthesizing cross-tenant connections.
@@
Two embedding models. We evaluate two open-source models (all-MiniLM-L6-v2, 384d, and all-mpnet-base-v2, 768d) and confirm the vulnerability persists across both (§VII-E). However, we do not test commercial embedding models (e.g., OpenAI text-embedding-3-large) which may produce different similarity distributions. The pivot vulnerability is structural (entity linking, not embedding quality), so we expect model independence to hold broadly.
@@
Adaptive attacker scope. We evaluate three adaptive attacks (A5–A7) including targeted metadata forgery at rates up to 10%, entity manipulation, and query manipulation (§VII-G). All defended pipelines maintain $\text{RPR}=0.0$. However, we do not evaluate attackers who craft high-provenance payloads to bypass D4 specifically, or adversarial query phrasing to manipulate D2’s query classifier. Our metadata forgery assumes the attacker can relabel node properties but cannot modify the graph schema or index structure.
NER and entity linker dependence. The pivot path exists because entity linking creates shared nodes across tenant boundaries. Both corpora use spaCy’s en_core_web_sm NER model, which has known recall limitations for domain-specific entities. On the Enron corpus, this produces noisier entity extractions (dates, monetary amounts, partial names) that create spurious cross-tenant connections—yet RPR is lower (0.695 vs. 0.954), suggesting that NER noise creates fragmented rather than dense pivot paths. A higher-recall linker would extract more shared entities, potentially increasing organic leakage. We do not evaluate how linker quality affects RPR, nor do we test production entity linking systems that perform coreference resolution or cross-document entity merging.
@@
XI Related Work
XI-A Vector RAG Poisoning

### v2

@@ 7.4 Connectivity Sensitivity
7.5 Embedding Model Sensitivity
To verify that the pivot vulnerability is structural rather than embedding-dependent, we repeat the P1 and P3 evaluations using all-mpnet-base-v2 (768 dimensions, higher retrieval quality) in addition to our default all-MiniLM-L6-v2 (384 dimensions). We rebuild the ChromaDB collection for each model and run the full 500-query evaluation (Table 6).
Table 6: Embedding model sensitivity: P3 RPR and mean Leakage@k across two embedding models. The vulnerability persists regardless of embedding quality.
@@ 7.5 Embedding Model Sensitivity
7.6 Attack Evaluation
Table 7 presents RPR under each attack across pipeline variants. All four attacks achieve $\text{RPR}=1.0$ against the undefended hybrid pipeline (P3)—every adversarial query produces cross-tenant leakage. Critically, all four attacks achieve $\text{RPR}=0.0$ against every defended variant (P4, P6, P8).
Table 7: RPR under each attack type across pipeline variants. Mean Leakage@k shown in parentheses. All attacks fail against D1-defended pipelines.
@@ 7.6 Attack Evaluation
Table 8 provides injection details for each attack. The attacks span a range of injection budgets (9–20 chunks) and entity strategies (1–2 target entities, 1–6 MENTIONS edges created).
Table 8: Attack injection details. All attacks achieve RPR=1.0 on P3 and RPR=0.0 on P4/P6/P8.
@@
7.7 Adaptive Attacks (A5–A7)
We extend the attack taxonomy with three adaptive strategies that target the defense mechanisms themselves (Table 9).
Table 9: RPR under adaptive attacks (A5–A7) across pipeline variants. A5 tests metadata forgery at three rates; A6 tests entity manipulation; A7 tests query manipulation. Leakage@k shown in parentheses.
@@ 7.7 Adaptive Attacks (A5–A7)
7.8 Generation Impact
To measure whether leaked context actually contaminates LLM-generated answers, we evaluate production LLMs on queries with known leakage across all three corpora (Table 10).
Table 10: Generation contamination metrics across datasets and LLMs. ECR = Entity Contamination Rate, ILS = Information Leakage Score, FCR = Factual Contamination Rate, GRR = Generation Refusal Rate. $n$ = queries with leakage evaluated.
Dataset LLM ECR ILS FCR GRR $n$
Synthetic GPT-4o 0.077 0.305 0.050 0.800 10
Claude Sonnet 4.5 0.321 0.352 0.072 0.047 64
DeepSeek-V3 0.147 0.333 0.044 0.045 134
Enron GPT-5.2 0.082 0.285 0.005 0.000 181
Claude Sonnet 4.5 0.133 0.280 0.004 0.000 40
DeepSeek-V3 0.053 0.260 0.056 0.050 101
EDGAR Claude Sonnet 4.5 0.040 0.358 0.000 0.019 53
DeepSeek-V3 0.025 0.345 0.016 0.101 139
Note: Synthetic Claude Sonnet 4.5 and EDGAR Claude Sonnet 4.5 used self-judging; all other rows used Claude Sonnet 4.5 as judge. Synthetic GPT-4o (n=10n{=}10) used self-judging. ECR drops from synthetic to Enron to EDGAR as entity distinctiveness decreases. EDGAR shows the lowest ECR ($\leq 0.04$) because financial entity names appear naturally in any financial answer.
The results reveal model-dependent and dataset-dependent contamination behavior. On the synthetic corpus, Claude Sonnet 4.5 exhibits the highest Entity Contamination Rate ($\text{ECR}=0.32$, n=64n{=}64), indicating it readily incorporates leaked entities into generated answers. GPT-4o shows the opposite pattern: $\text{ECR}=0.08$ and $\text{GRR}=0.80$ (n=10n{=}10), largely ignoring leaked context. DeepSeek-V3 falls between ($\text{ECR}=0.15$, n=134n{=}134) with low GRR ($0.05$) and FCR ($0.04$). All three models show GRR $\leq 0.05$ at scale, suggesting that generation refusal is rare when leaked context is topically relevant.
On the Enron corpus (n=181n{=}181 for GPT-5.2, n=40n{=}40 for Claude, n=101n{=}101 for DeepSeek), ECR drops substantially across all models (GPT-5.2 $0.08$; Claude $0.13$; DeepSeek $0.05$) while GRR approaches zero (all models $\leq 0.05$). Real email content is contextually relevant enough that LLMs rarely ignore it, but leaked entities from adjacent departments are less distinctive than synthetic ones, producing lower entity contamination. GPT-5.2 shows moderate ECR ($0.082$) with near-zero FCR ($0.005$), suggesting it surfaces entity names but avoids reproducing substantive facts from leaked context. These findings confirm that retrieval-level leakage translates to generation-level contamination across all three corpora, that the severity depends on the model and the domain, and that the effect is robust across sample sizes (n=40n=40181181).
7.9 Traversal Parameter Sweep
@@
7.10 Latency and Overhead
Table 11 presents latency and context size across pipeline variants.
Table 11: Latency (ms) and context size across pipeline variants.
@@ 7.10 Latency and Overhead
7.11 Metadata Integrity Stress Test
@@ 7.13 Utility Impact
Table 12 presents the security-utility tradeoff. P3 (undefended hybrid) returns 110 items per query, but only 85.5% are authorized—the remaining 16 items (14.5%) are cross-tenant or over-clearance leakage. P1 (vector-only) returns 10 fully authorized items, establishing the vector-only utility baseline.
@@
7.14 Cross-Dataset Validation
To verify that Retrieval Pivot Risk is not an artifact of synthetic corpus construction, we repeat the baseline evaluation (P1, P3, P4) on the Enron email corpus (§6.3) and EDGAR 10-K filings (§6.4). Table 13 compares results across all three corpora.
Table 13: Cross-dataset baseline: retrieval security metrics for undefended (P3) and defended (P4) hybrid RAG across three corpora. RPR = Retrieval Pivot Risk, Leak = mean leaked items per query, PD = Pivot Depth (hops), Ctx = mean context size.
@@ 7.14 Cross-Dataset Validation
EDGAR P1 (Vector) 0.000 0.0 10
P3 (Hybrid) 0.085 0.4 2.0 39
P4 (+D1) 0.000 0.0 22
Note: Synthetic corpus: 1,000 documents, 4 tenants, 40 bridge entities. Enron corpus: 50,000 emails, 5 departments, 19 cross-department entities. EDGAR corpus: 887 10-K filing sections, 4 sector-tenants, 13 cross-sector entities. All corpora evaluated with 200 benign queries.
Four findings emerge from the cross-dataset comparison:
The vulnerability generalizes. Enron’s undefended hybrid pipeline (P3) produces $\text{RPR}=0.695$ on benign queries—139 of 200 queries leak cross-department content, with a mean of 7.1 leaked items per query. EDGAR produces $\text{RPR}=0.085$—lower but still affecting 17 of 200 queries. The vulnerability is not an artifact of synthetic construction: real corporate data with natural organizational boundaries exhibits the same pivot mechanism.
RPR scales with entity connectivity. The three corpora span a wide range: synthetic ($\text{RPR}=0.954$, 40 bridge entities), Enron ($\text{RPR}=0.695$, 19 shared entities), and EDGAR ($\text{RPR}=0.085$, 13 cross-sector entities). RPR correlates with the number and density of cross-tenant entity connections. The synthetic graph has curated cross-references, Enron has informal cross-department email traffic, and EDGAR’s formal financial language creates fewer organic entity connections across sectors. This confirms that RPR is a function of entity connectivity density, not corpus size.
D1 eliminates all leakage on all three corpora. P4 achieves $\text{RPR}=0.0$ on Enron (context size 25), EDGAR (context size 22), and synthetic (context size 56). The defense generalizes without modification: the same per-hop tenant-and-sensitivity check eliminates leakage regardless of corpus size, entity distribution, or graph topology.
PD = 2 persists. All leakage in all three corpora occurs at exactly $\text{PD}=2$, confirming that the bipartite pivot (chunk $\to$ entity $\to$ chunk) is a structural invariant of hybrid RAG systems that construct knowledge graphs via entity linking, independent of the underlying document domain.
8 Mitigations
We propose five defenses that operate at different stages of the hybrid pipeline, forming a defense-in-depth architecture (Table 14).
Table 14: Defense mechanisms, pipeline integration points, and experimental impact.
@@ 8.1 D1: Per-Hop Authorization
This entity-level filtering is precisely the mechanism by which D1 eliminates cross-tenant leakage. The 2-hop pivot path (chunk $\to$ entity $\to$ chunk) requires traversal through a shared entity node. By filtering entity nodes, D1 severs this path at hop 1, preventing the traversal from reaching the unauthorized chunk at hop 2. The trade-off is that entity information (which may be useful for answer generation) is excluded from the final context. We quantify this impact in Section 9.
@@
8.2 D1 Metadata Integrity Assumption
D1 assumes that node metadata (tenant, sensitivity) is trustworthy. If an attacker can forge metadata during document injection, D1 can be bypassed. Our metadata stress test (§7.11) shows that D1 is robust to random mislabeling up to 5%, but targeted metadata forgery is a different threat. Mitigations include: (a) system-assigned metadata during ingestion (the uploader cannot choose tenant or sensitivity labels), (b) cryptographic signing of metadata at ingestion time, and (c) periodic metadata audits that compare node labels against source-of-truth identity systems. In practice, most enterprise document management systems already enforce system-assigned tenant labels, making (a) the natural deployment model.
8.3 D2: Edge Allowlist
@@
8.4 D3: Budgeted Traversal
Budgeted traversal enforces hard limits on BFS expansion: maximum hop depth, maximum branching factor per node, and maximum total expanded nodes. Our traversal sweep (§7.9) shows that $\text{total\_nodes}\leq 25$ eliminates leakage even without D1, while the default budget ($d_{\text{max}}=2$, branching $\leq 10$, total $\leq 50$) reduces context from 56 items (D1 only) to 28 items, removing authorized but irrelevant graph content that adds noise.
8.5 D4: Trust-Weighted Expansion
@@
8.6 D5: Merge-Time Policy Filter
@@
9 Discussion
9.1 Security-Utility Tradeoff
@@
D1’s security guarantee comes at a specific cost: entity nodes are excluded from the final context. In our bipartite construction, the undefended P3 pipeline returns approximately 110 items per query, of which roughly 25–35 are entity nodes. After D1 filtering, these entity nodes are removed entirely, leaving 50–56 chunk-only items. The entity information they carried (names, relationships, organizational context) is lost. For queries that benefit from entity-level knowledge (e.g., “Who works on project Alpha?”), this may reduce answer quality compared to an oracle that retains only authorized entity nodes. We consider this an acceptable trade-off: the entity nodes are exactly the pivot vectors that enable cross-tenant leakage, and their removal is the mechanism by which D1 achieves $\text{RPR}=0.0$. Future work on multi-tenant entity authorization (Section 10) could recover entity utility while preserving security.
@@
9.2 Why PD = 2 Across All Three Corpora
@@
The Enron and EDGAR corpora confirm this structural invariant: despite graphs ranging from 2.8K to 376K nodes, varying NER quality, and document types from curated templates to informal email to formal financial filings, all leakage occurs at exactly $\text{PD}=2$. The synthetic corpus has 40 shared entities, Enron has 19, and EDGAR has 13. All three produce the same pivot depth. This confirms that $\text{PD}=2$ is a consequence of the bipartite construction, not of corpus-specific properties.
@@ 9.3 The Amplification Mechanics
9.4 Comparison with Prior Art
@@
GRAGPoison [gragpoison2026] demonstrates that graph structure amplifies small poisoning inputs (98% ASR from relation injection). Our work extends this insight to hybrid architectures: the combination of vector retrieval and graph expansion creates a compound attack surface where authorized vector results seed unauthorized graph traversals.
@@
PoisonedRAG [zou2025poisonedrag] achieves 90% ASR by injecting 5 documents into vector stores. In vector-only RAG, this is contained by prefilters. In hybrid RAG, even a single retrieved document that mentions a shared entity can trigger graph expansion into sensitive neighborhoods, making the effective attack surface orders of magnitude larger.
@@
The Graph RAG Privacy Paradox [graphragprivacy2025] observes that graph RAG reduces text leakage but increases structural information extraction. Our findings align: the hybrid pipeline’s graph expansion systematically transforms authorized text into unauthorized structural access.
9.5 Implications for Agentic Systems
@@
The pivot attack is especially dangerous in agentic RAG deployments (LangGraph, CrewAI) where graph traversal is performed autonomously by an LLM agent [multiagentic2025]. In these systems, the agent decides traversal depth, edge types, and expansion strategies without human oversight. A poisoned vector seed can manipulate the agent into performing unrestricted graph exploration, and research shows that a single compromised agent can poison 87% of downstream decision-making within 4 hours [multiagentic2025]. Per-hop authorization is even more critical in agentic settings, where there is no human in the loop at the expansion step.
10 Limitations
@@
Three corpora. We evaluate on a synthetic enterprise corpus, the Enron email corpus, and SEC EDGAR 10-K filings. The synthetic corpus validates the vulnerability under controlled conditions; the Enron and EDGAR corpora confirm generalization to real-world data with varying entity connectivity densities. However, all three use a bipartite chunk–entity graph topology that produces a uniform PD=2 signature. Knowledge graphs with richer entity-to-entity relationships (ontological hierarchies, multi-hop inference chains) may exhibit leakage at deeper pivot depths. We do not evaluate on production enterprise graphs with RBAC/ABAC policies, which may exhibit different connectivity patterns.
@@
Generation evaluation scope. Our generation contamination experiment (§7.8) evaluates three LLMs across all three corpora (n=40n{=}40–181 leaking queries on Enron, n=64n{=}64–134 on synthetic, n=53n{=}53–139 on EDGAR) and confirms that leaked context contaminates generated answers (ECR up to 0.32, FCR up to 0.07). The GPT-4o synthetic sample (n=10n{=}10) predates our expanded query pool; larger samples across Enron and EDGAR provide more robust estimates. We do not evaluate chain-of-thought reasoning, which might amplify the impact by synthesizing cross-tenant connections.
@@
Two embedding models. We evaluate two open-source models (all-MiniLM-L6-v2, 384d, and all-mpnet-base-v2, 768d) and confirm the vulnerability persists across both (§7.5). However, we do not test commercial embedding models (e.g., OpenAI text-embedding-3-large) which may produce different similarity distributions. The pivot vulnerability is structural (entity linking, not embedding quality), so we expect model independence to hold broadly.
@@
Adaptive attacker scope. We evaluate three adaptive attacks (A5–A7) including targeted metadata forgery at rates up to 10%, entity manipulation, and query manipulation (§7.7). All defended pipelines maintain $\text{RPR}=0.0$. However, we do not evaluate attackers who craft high-provenance payloads to bypass D4 specifically, or adversarial query phrasing to manipulate D2’s query classifier. Our metadata forgery assumes the attacker can relabel node properties but cannot modify the graph schema or index structure.
NER and entity linker dependence. The pivot path exists because entity linking creates shared nodes across tenant boundaries. All three corpora use spaCy’s en_core_web_sm NER model, which has known recall limitations for domain-specific entities. On the Enron corpus, this produces noisier entity extractions (dates, monetary amounts, partial names) that create spurious cross-tenant connections—yet RPR is lower (0.695 vs. 0.954), suggesting that NER noise creates fragmented rather than dense pivot paths. A higher-recall linker would extract more shared entities, potentially increasing organic leakage. We do not evaluate how linker quality affects RPR, nor do we test production entity linking systems that perform coreference resolution or cross-document entity merging.
@@
11 Related Work
11.1 Vector RAG Poisoning